# Chiral anomaly-induced nonlinear Hall effect in spin-orbit coupled noncentrosymmetric metals

### v2

@@ II.2.1 Changes in Maxwell-Boltzmann transport (MBT) theory due to shift of origin
where the domains of integration are defined over iso-energetic surfaces in momentum space as: $\mathcal{D}_{\bm{\mathcal{K}}}=\left\{\mathcal{K}(\theta,\phi)\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ \mathcal{K}(\theta,\phi)=\mathcal{K}_{\mathrm{F}},\leavevmode\nobreak\ \mathcal{E}(\mathcal{K}(\theta,\phi))=\mathcal{E}_{\mathrm{F}}\right\},\quad\mathcal{Q}_{\bm{\mathcal{Q}}}=\left\{\mathcal{Q}(\theta,\phi)\leavevmode\nobreak\ \big{|}\leavevmode\nobreak\ \mathcal{Q}(\theta,\phi)=\mathcal{Q}_{\mathrm{F}},\leavevmode\nobreak\ \mathcal{E}(\mathcal{Q}(\theta,\phi))=\mathcal{E}_{\mathrm{F}}\right\}$. These correspond to constant energy contours defined on the Fermi surface. The dispersion relation derived from Eq. 23 yields isoenergy manifolds that are analogous to those reported in Ref. [35] for spin-orbit coupled noncentrosymmetric Weyl metals (SOC-NCMs), thereby enabling a direct comparison of their topological and transport features.
@@
II.2.2 Berry curvature for a general two-band Hamiltonian.
@@
Here, $\mathbf{k}$ is wave vector and $\bm{\sigma}$ is vector of Pauli matrices. $\bm{\mathcal{W}}(\mathbf{k})=\mathcal{W}\leavevmode\nobreak\ (\sin{\alpha}\cos{\beta},\leavevmode\nobreak\ \sin{\alpha}\sin{\beta},\leavevmode\nobreak\ \cos{\alpha})$ with $0\leq\alpha\leq\pi$, $0\leq\beta\leq 2\pi$. Then, the Eq. 27 can be written as,
$H_{2\times 2}(\mathbf{k})=\mathcal{W}\leavevmode\nobreak\ \begin{bmatrix}\cos{\alpha}&\sin{\alpha}\leavevmode\nobreak\ e^{-i\beta}\\ \sin{\alpha}\leavevmode\nobreak\ e^{i\beta}&-\cos{\alpha}\end{bmatrix}$ (28)
@@
Figure 6: A schematic illustration depicting the characteristic regimes of magnetoconductivity $\sigma_{ij}$ in chiral Weyl semimetals, highlighting the phenomena of weak sign reversal, strong sign reversal, and the coexisting regime of strong-and-weak sign reversal, in contrast to the conventional quadratic-in-$B$ magnetoconductivity response [51, 35].
$\Omega_{xy}=-2\leavevmode\nobreak\ \mathrm{Im}\biggr{\langle}\frac{\partial u}{\partial k_{x}}\biggr{|}\frac{\partial u}{\partial k_{x}}\biggr{\rangle}.$ (29)
@@
III Results
Before discussing the CNLH, here we would like to briefly review the concept of different types of sign reversal. The sign of longitudinal magnetoconductivity (LMC) in Weyl materials has been the focus of extensive study in previous research. In the case of untilted Weyl semimetals, the chiral anomaly is expected to produce a positive LMC under conditions of weak interband scattering; however, with sufficiently strong interband scattering, the sign of LMC reverses [57, 53, 49]. Conversely, even a slight tilting of the Weyl cone can lead to negative LMC along a specific direction of the magnetic field, even in the regime of weak interband scattering. Despite the similarity in sign reversal, the underlying mechanisms in these scenarios are fundamentally distinct, giving rise to the classifications of ‘strong-sign-reversal’ and ‘weak-sign-reversal,’ as outlined in Ref. [51, 35]. We provide a brief review here. A general expression for the magnetoconductivity tensor can be expressed as [51]
$\sigma_{ij}(B)=\sigma_{ij}^{(0)}+(B-B_{0})^{2}\sigma_{ij}^{(2)},$ (31)
which integrates (i) normal quadratic $B$-dependence, (ii) linear-in-$B$ dependence with sign change along a specific magnetic field direction, and (iii) quadratic-in-$B$ dependence with a negative sign, within a unified framework. The features distinguishing ‘weak-sign-reversal’ include (i) $B_{0}\neq 0$, (ii) $\sigma_{ij}^{(0)}\neq\sigma_{ij}(B=0)$, and (iii) $\mathrm{sign}\;\sigma_{ij}^{(2)}>0$. In this situation, the vertex of the magnetoconductivity parabola shifts away from the origin, and the conductivity exhibits opposite signs for small positive and negative magnetic fields. However, the parabola’s orientation remains positive, i.e., $\mathrm{sign}\;(\sigma_{ij}^{(2)})>0$. In contrast, ‘strong-sign-reversal’ is identified by $\mathrm{sign}\;(\sigma_{ij}^{(2)})<0$, which signifies a complete inversion of the parabola’s orientation. Tilting of Weyl cones can result in ‘weak-sign-reversal,’ while strong interband scattering or strain typically leads to ‘strong-sign-reversal’ [51, 35]. Remaining possibility, i.e., completely reverse and shifted parabola results in ‘strong and weak-sign-reversal’. Fig. 6 illustrates the distinction between these three cases schematically.
Figure 7: The chiral anomaly-induced nonlinear Hall response (CNLH) is analyzed as a function of the tilt parameter $\alpha$ and the angle $\gamma$, at a fixed magnetic field strength of $B=0.50\leavevmode\nobreak\ \mathrm{T}$. Panel (a) corresponds to the case of point-like non-magnetic impurity scattering, while panels (b), (c), and (d) represent magnetic impurity scattering with spin orientation aligned along the $\hat{x}-$, $\hat{y}-$, and $\hat{z}-$directions, respectively. The omission of the orbital magnetic moment (OMM) in our analysis affects only the overall amplitude—typically leading to a reduction—without altering the qualitative behavior. Hence, OMM has been included in all the present discussion.
III.1 CNLH in absence of tilt
In the absence of tilt in the band dispersion in SOC-NCMs, the chiral anomaly-induced nonlinear Hall (CNLH) conductance exhibits a quadratic dependence on the magnetic field $B$ and remains symmetric about $B=0$. In contrast to the linear-response regime—where a positive longitudinal magnetoconductance is typically observed due to the chiral anomaly—we consistently observe a negative magnetoconductance in the nonlinear regime. This behavior, characterized by a sign opposite to the linear response, is referred to here as a strong sign reversal. To examine the robustness of this feature, we analyze the behavior of CNLH conductance under varying strengths of interband scattering, denoted by $\alpha$. As expected, the magnitude of negative conductance decreases (i.e., becomes less negative) with increasing $\alpha$, indicating a suppression of anomaly-induced transport due to enhanced decoherence between bands. As established in earlier studies, the dependence of the conductance on the angle $\gamma$ between the magnetic field $\mathbf{B}$ and increasing $x$-direction shows that CNLH vanishes at $\gamma=0$ and $\gamma=\pi/2$, reaching a maximum at $\gamma=\pi/4$. The vanishing of CNLH at $\gamma=\pi/2$ is attributed to azimuthal symmetry, which nullifies the $\phi$-dependent integrand contributions. At $\gamma=0$, the projection $\mathbf{E}\cdot\mathbf{B}$ associated with the chiral anomaly becomes zero, thereby eliminating the anomaly-induced transport response. Furthermore, we investigate the influence of impurity type on the magnitude and profile of the CNLH conductance. We consider four distinct impurity scattering profiles: non-magnetic ($\sigma_{0}$) and magnetic impurities polarized along orthogonal directions ($\sigma_{x}$, $\sigma_{y}$, and $\sigma_{z}$). Although the overall qualitative behavior of the CNLH signal remains consistent across these impurity types, we observe significant quantitative variations. In particular, magnetic impurity types $\sigma_{x}$ and $\sigma_{y}$ lead to a larger magnitude of conductance, as illustrated in Fig. 7, whereas $\sigma_{0}$ and $\sigma_{z}$ impurities yield comparatively smaller conductance amplitudes.
Figure 8: The chiral anomaly-induced nonlinear Hall conductivity (CNLH) is analyzed as a function of the dimensionless interband scattering strength $(\alpha)$, for various values of the tilt parameter $t_{z}$, under a fixed magnetic field $B=0.50\leavevmode\nobreak\ \text{T}$ oriented along the direction $\gamma=\pi/4$. The arrows indicate the progression of $t_{z}$ from $0.98\,v_{F}$ to $0.98\,v_{F}$, illustrating the effect of tilt modulation on the nonlinear transport response.
Figure 9: The chiral anomaly-induced nonlinear Hall response (CNLH) is analyzed as a function of tilt along the $z$-axis ($t_{z}$) for various values of interband scattering strength ($\alpha$), under a fixed magnetic field $B=0.50\leavevmode\nobreak\ \text{T}$ oriented at an angle $\gamma=\pi/4$. Similar to the case with finite $t_{x}$, the CNLH vanishes at a finite, nonzero value of the tilt parameter $t_{z}$, indicating a symmetry-breaking threshold. For all four impurity scattering configurations considered—encompassing scalar and magnetic disorder profiles—the CNLH exhibits an approximately linear and monotonic dependence on $t_{z}$ within the parameter regime explored.
III.2 CNLH in the presence of tilt along the z-direction
III.3 CNLH in the presence of tilt along the x-direction
Figure 10: The chiral anomaly-induced nonlinear Hall conductivity (CNLH) is plotted as a function of the dimensionless interband scattering strength parameter ($\alpha$) for various values of the tilt parameter ($t_{x}$), under a fixed magnetic field $B=0.50\leavevmode\nobreak\ \text{T}$ applied at an oblique angle characterized by $\gamma=\pi/4$ with respect to $\hat{x}$. The dashed curves correspond to negative tilt configurations, i.e., $t_{x}/v_{F}<0$. Panels (a) and (d) exhibit qualitatively similar CNLH responses, whereas panels (b) and (c) reveal distinct deviations due to the interplay between the tilt-induced anisotropy and interband scattering.
Figure 11: The chiral anomaly-induced nonlinear Hall (CNLH) conductivity is plotted as a function of the tilt parameter ($t_{x}$) for various interband scattering strengths ($\alpha$), under a fixed magnetic field magnitude $B=0.50\leavevmode\nobreak\ \mathrm{T}$ oriented along the direction $\gamma=\pi/4$. Notably, the CNLH response vanishes at a finite (non-zero) value of tilt, indicating a nontrivial cancellation mechanism. With the exception of panel (b), the CNLH exhibits a predominantly linear and monotonic dependence on $t_{x}$, reflecting the symmetry-breaking role of tilt in modulating the anomalous transport.
IV Conclusion
V Acknowledgments

### v3

@@ Appendix B Changes in Maxwell-Boltzmann transport (MBT) theory due to shift of origin
where, $\mathcal{F}_{1}(\mathbf{\bm{\mathcal{K}}})$ and $\mathcal{F}_{2}(\mathbf{\bm{\mathcal{Q}}})$ generic functions of $\mathbf{\bm{\mathcal{K}}}$ and $\mathbf{\bm{\mathcal{Q}}}$. The domains of integration are defined over iso-energetic surfaces in momentum space as: $\mathcal{D}_{\bm{\mathcal{K}}}=\left\{\mathcal{K}(\theta,\phi)\penalty 10000\ \big|\penalty 10000\ \mathcal{K}(\theta,\phi)=\mathcal{K}_{\mathrm{F}},\penalty 10000\ \mathcal{E}(\mathcal{K}(\theta,\phi))=\mathcal{E}_{\mathrm{F}}\right\},\quad\mathcal{Q}_{\bm{\mathcal{Q}}}=\left\{\mathcal{Q}(\theta,\phi)\penalty 10000\ \big|\penalty 10000\ \mathcal{Q}(\theta,\phi)=\mathcal{Q}_{\mathrm{F}},\penalty 10000\ \mathcal{E}(\mathcal{Q}(\theta,\phi))=\mathcal{E}_{\mathrm{F}}\right\}$. These correspond to constant energy contours defined on the Fermi surface. The dispersion relation derived from Eq. 3 yields isoenergy manifolds that are analogous to those reported in Ref. [varma2024magnetotransport] for spin-orbit coupled noncentrosymmetric Weyl metals (SOC-NCMs), thereby enabling a direct comparison of their topological and transport features.
@@
Appendix C Berry curvature for a general two-band Hamiltonian.
@@
Here, $\mathbf{k}$ is wave vector and $\bm{\sigma}$ is vector of Pauli matrices. $\bm{\mathcal{W}}(\mathbf{k})=\mathcal{W}\penalty 10000\ (\sin{\alpha}\cos{\beta},\penalty 10000\ \sin{\alpha}\sin{\beta},\penalty 10000\ \cos{\alpha})$ with $0\leq\alpha\leq\pi$, $0\leq\beta\leq 2\pi$. Then, the Eq. 30 can be written as,
$H_{2\times 2}(\mathbf{k})=\mathcal{W}\penalty 10000\ \begin{bmatrix}\cos{\alpha}&\sin{\alpha}\penalty 10000\ e^{-i\beta}\\ \sin{\alpha}\penalty 10000\ e^{i\beta}&-\cos{\alpha}\end{bmatrix}$ (31)
@@
$\Omega_{xy}=-2\penalty 10000\ \mathrm{Im}\biggr\langle\frac{\partial u}{\partial k_{x}}\biggr|\frac{\partial u}{\partial k_{x}}\biggr\rangle.$ (32)